OpenStudy (anonymous):

4x-2y=4 x=4 what is y

OpenStudy (anonymous):

y=-6

OpenStudy (anonymous):

4(4)-2y=4 16-2y=4 -2y=4-16 -2y=-12 y=-12/12 y=6

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

nevermind you said -6 and it was just 6

OpenStudy (anonymous):

sorry for the minus. was typing mistake

OpenStudy (anonymous):

its ok thank you